# GDSR: Global-Detail Integration through Dual-Branch Network with Wavelet Losses for Remote Sensing Image Super-Resolution

### v1

@@ IV Experiment
IV-A Datasets and Evaluation
This paper reports the performance of SR on two remote sensing image datasets, including the publicly available AID dataset [39] and the custom RSSRD-QH dataset.
The AID dataset consists of 30 categories. For each category, the data was randomly split into two equal parts, one for training and the other for testing. Additionally, 20% of the training data was further separated and used as a validation set. Specifically, the training set contained 4,000 images, the validation set 1,000 images, and the test set 5,000 images.
The RSSRD-QH dataset is located in representative watershed units within the Sanjiangyuan region and the Qinghai Lake surrounding area. The geographical distribution of the counties covered by the dataset, including their longitude and latitude ranges, is summarized in Table I. The dataset was collected using unmanned aerial vehicles, with flight altitudes carefully controlled between 30 and 50 meters. The average spatial resolution of the captured imagery is 0.01 meters, providing comprehensive coverage of various landscape elements across different elevation gradients, including terraces, floodplains, gentle slopes, and steep slopes within these watershed units. From each Watershed, we randomly selected 50% of the images for the training set and 50% for the test set, with 20% of the training set randomly selected for validation. This resulted in a training set of 2,242 images, a validation set of 561 images, and a test set of 2,804 images. The HR images in both datasets were 600 × 600 pixels, and the corresponding LR images were 200 × 200 pixels.
The proposed method, along with other competing methods, was evaluated on the test datasets using two classic full-reference metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [40]. Additionally, the LPIPS [41] metric was used to capture the perceptual quality of the images. Notably, PSNR and SSIM were calculated based on the luminance channel (Y) of the YCbCr color space, while LPIPS evaluated the perceptual similarity between images by extracting features from a pre-trained deep network, specifically the AlexNet model [42], which was employed in this study.
IV-B Degradation Model and Implementation Details
We employed both bicubic downsampling and a comprehensive degradation model (CDM) to simulate LR images during training. Based on existing blind SR methods [43, 44] and the characteristics of remote sensing sensors [45], the CDM for LR image synthesis included isotropic Gaussian blur, motion blur, scaling with different interpolation methods, additive Gaussian noise, and JPEG compression noise, all applied in random order.
Our GDSR performed deep feature exploration through 4 RGEGs, RDEGs, and GDRMs. Each RGEG consisted of 6 R-RWKVBs and each RDEG contains 12 RCBs. Empirically, we set the internal channel dimension to = 96. For the GDSR_TC and GDSR_MC, the RGEG branches correspond to the RTSB and ResidualGroup of SwinIR [14] and MambaIR [18], respectively, with all other settings remaining consistent with GDSR.
During training, each HR image was randomly cropped to a size of 192 × 192 pixels, with the corresponding LR image size being 64 × 64 pixels. The batch size was set to 16, and the Adam optimizer was used with = 0.9 and = 0.999. The initial learning rate was set to , which was halved at the 100th epoch, with a total of 200 training epochs. All SR models were implemented in the PyTorch framework. The FLOPs results were calculated using an input tensor of size 1 × 3 × 240 × 240, and the computations were performed using torch-operation-counter [46]. The inference times were tested on 300 random images with a size of 3 × 240 × 240. All experiments were conducted on an NVIDIA A800 80GB GPU.
RCBs | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
1 | 28.96/0.7658 | 0.3514 | 5.86M | 340.25G | 7.3 |
4 | 28.96/0.7657 | 0.3518 | 7.85M | 455.31G | 7.0 |
8 | 29.00/0.7677 | 0.3499 | 10.51M | 608.73G | 6.7 |
12 | 29.03/0.7686 | 0.3476 | 13.17M | 762.14G | 6.4 |
16 | 29.01/0.7680 | 0.3480 | 15.83M | 915.55G | 6.1 |
20 | 29.03/0.7688 | 0.3477 | 18.48M | 1068.97G | 5.8 |
@@ IV-C Ablation Study
In this section, we discuss the proposed GDSR in depth by investigating the effect of its major components and their variants.
@@ IV-C1 Effect of Key Modules
(a) Effect of RDEG. To investigate the effect of RDEG on the model performance, we conducted ablation experiments by varying the number of RCBs in RDEG.
The experimental results, presented in Tables II and III, indicated that the model achieved a PSNR of 28.94 dB and an SSIM of 0.7652 when only the RGEG branch was utilized without incorporating RDEG. Adding a single RCB significantly improved the model’s performance, increasing the PSNR to 28.96 dB and the SSIM to 0.7658. This improvement highlighted the RCB’s capability to capture critical information, thereby enhancing reconstruction quality.
As the number of RCBs increased, further improvements in PSNR and SSIM were observed. For instance, increasing the RCB count from 1 to 20 raised the PSNR from 28.96 dB to 29.03 dB and the SSIM from 0.7658 to 0.7688. However, the rate of improvement diminished as the RCB count increased, and beyond a certain threshold, performance even began to decline. This trend suggested that while adding more RCBs initially benefited the model’s output, the marginal gains decreased and eventually plateaued. Additionally, increasing the number of RCBs led to a linear rise in computational complexity, as evidenced by the increase in parameters, FLOPs, and a decrease in FPS. This highlighted the trade-off between performance gains and computational efficiency. Therefore, determining the optimal number of RCBs is crucial in practical applications, especially in scenarios where computational efficiency is a critical factor.
(b) Effect of GDRM. To evaluate the effectiveness of the GDRM, we conducted ablation experiments separately using the RDEG and RGEG branches. The experimental results were presented in Table III. The comparisons demonstrated that introducing the GDRM significantly improved performance compared to single-branch models. The GDSR model achieved superior PSNR, SSIM, and LPIPS values, indicating that GDSR effectively integrated the local detail information from RDEG and the global contextual information from RGEG. Additionally, the incorporation of PSAM enabled GDRM to fuse the information from both branches more effectively.
(c) Visualization of GDRM. Fig. 7 visualized the features of the RDEG, RGEG branches, and GDRM. We observed that the feature maps from RDEG and RGEG were complementary, while GDRM combined features from both branches. This demonstrated that, within the dual-branch structure of GDSR, each branch focused on different features, and fusing these features enabled effective image reconstruction.

### v4

@@ IV-C Ablation Study
In this section, we discuss the proposed GDSR in depth by investigating the effect of its major components and their variants. All experiments were conducted under the ×3 super-resolution configuration with the loss function in the AID [44] dataset.
@@ IV-C1 Effect of Key Modules
(a) Effect of RDEG and RGEG. Ablation studies were conducted to evaluate the impact of varying RCBs in RDEG and R-RWKVBs in RGEG on model performance. As shown in Tables II and III, the baseline model using only the RGEG branch yielded a PSNR of 28.94 dB and SSIM of 0.7652. Introducing a single RCB to RDEG improved performance to 28.96 dB PSNR and 0.7658 SSIM, validating the RCB’s ability to enhance reconstruction quality. Scaling the RCB count to 12 further increased PSNR to 29.03 dB and SSIM to 0.7686, with a computational footprint of 13.17M parameters and 338.73G FLOPs. However, beyond 12 RCBs, performance plateaued (e.g., 29.01 dB with 16 RCBs) or exhibited instability (e.g., 29.03 dB with 20 RCBs), while computational costs grew linearly, reducing FPS from 14.8 to 12.6. This suggests diminishing returns from excessive RCB stacking, necessitating a balance between accuracy and efficiency.
Similarly, experiments on RGEG’s R-RWKVBs revealed analogous trends. Fixing RCBs at 12 and increasing R-RWKVBs from 1 to 8 marginally improved PSNR and SSIM but incurred significant computational overhead, with parameters rising from 10.62M to 14.19M and FPS dropping sharply from 40.9 to 10.5. Notably, the joint configuration of 12 RCBs and 6 R-RWKVBs achieved optimal trade-offs, outperforming isolated branch evaluations (as shown in Table III). These results underscore the complementary roles of RDEG and RGEG in feature refinement while emphasizing the necessity of constrained component counts to mitigate redundancy and computational bloat in practical deployments.
| RCBs | R-RWKVBs | PSNR | SSIM | #Param. | FLOPs | FPS | |
| 1 |
28.96 | 0.7658 | 5.86M | 151.23G | 15.7 | |
| 4 | 28.96 | 0.7657 | 7.85M | 202.37G | 15.2 | ||
| 8 | 29.00 | 0.7677 | 10.51M | 270.55G | 14.5 | ||
| 12 | 29.03 | 0.7686 | 13.17M | 338.73G | 14.8 | ||
| 16 | 29.01 | 0.7680 | 15.83M | 406.92G | 13.3 | ||
| 20 | 29.03 | 0.7688 | 18.48M | 475.10G | 12.6 | ||
1 | 29.00 | 0.7676 | 10.62M | 272.97G | 40.9 | |
| 2 | 29.01 | 0.7678 | 11.13M | 286.12G | 29.4 | ||
| 4 | 29.01 | 0.7679 | 12.15M | 312.43G | 16.2 | ||
| 8 | 29.01 | 0.7679 | 14.19M | 365.04G | 10.5 |
(b) Effect of GDRM. Ablation experiments in Table III demonstrated that the dual-branch GDSR model outperformed single-branch alternatives, achieving superior PSNR, SSIM, and LPIPS values. This performance gain was attributed to the GDRM, which systematically integrated dual-branch features through its PSAM and adaptive fusion mechanism (Eqs. 14 and 15). From a spatial-frequency duality perspective, RGEG’s global attention mechanisms implicitly exhibited low-pass filtering characteristics to maintain structural continuity, while RDEG’s convolutional operations inherently emphasized high-frequency components for local detail enhancement [21]. As visualized in Fig. 7, the dual-branch structure of GDSR demonstrated complementary feature specialization with GDRM dynamically fusing both to enable efficient reconstruction.
Visualizations in Fig. 8 revealed that RGEG activated distributed pixels across broader regions, whereas RDEG focused on local patterns. The F2D strategy further enhanced this synergy by routing refined features exclusively to RDEG, prioritizing localized refinement over redundant global propagation. This design addressed a critical theoretical limitation: while RGEG aggregated broader pixel information, effective reconstruction required targeted utilization of spatial relationships within receptive fields. F2D’s spectral separation strategy specifically resolved conflicts observed in F2B, where dual-path feedback introduced frequency interference, as global attention diluted high-frequency details while convolutions struggled to recover them.
The quantitative results confirmed that GDSR’s dual-branch framework, coupled with GDRM’s reconstruction, systematically unified heterogeneous features to achieve both perceptual and pixel-wise accuracy. This validated the theoretical rationale that explicitly decoupling global (low-frequency) and local (high-frequency) processing paths, followed by attention-driven fusion, was essential for overcoming the limitations of conventional single-path architectures.
@@
(c) Effect of RWKV. To investigate the effect of RWKV, ablation experiments were conducted by replacing RWKV with Transformer [12] (GDSR_TC) and Mamba [15] (GDSR_MC) while keeping other components unchanged. As shown in Table IV, the RWKV-based GDSR achieves the best performance in terms of PSNR, SSIM, and LPIPS, outperforming both Transformer and Mamba variants. Notably, RWKV demonstrates superior computational efficiency with 14.8 FPS, significantly exceeding GDSR_TC and GDSR_MC. While maintaining moderate parameter size and FLOPs, RWKV achieves an optimal balance between reconstruction quality and computational cost. This validates RWKV’s effectiveness in capturing long-range dependencies while mitigating the quadratic complexity limitations of conventional attention mechanisms.
(d) Effect of Network Depth on Performance. To investigate the effect of network depth on the performance of GDSR, we varied the shared depth of both the RDEG and RGEG branches while keeping other design parameters constant. As illustrated in Table V, the experimental outcomes consistently demonstrated an enhancement in performance with deeper networks. Moreover, an evaluation of the trade-offs in terms of computational complexity was conducted. As the network depth increased, the model’s parameter count, FLOPS, and FPS were also scaled accordingly. These results supported the effectiveness of GDSR’s dual-branch design, where the shared depth allowed for both high-quality super-resolution and manageable computational overhead.
| Method | PSNR/SSIM | LPIPS | #Param. | FLOPs | FPS |
| GDSR_TC | 28.89/0.7633 | 0.3565 | 12.83M | 337.26G | 7.6 |
| GDSR_MC | 28.97/0.7663 | 0.3514 | 13.99M | 350.35G | 9.1 |
| GDSR | 29.03/0.7686 | 0.3476 | 13.17M | 338.73G | 14.8 |